OpenStudy (anonymous):

what is the solution to the system of equations below? 2x+3y=17 and 3x+6y=30

OpenStudy (anonymous):

well since it is a system of equations you need to combine it and cancel out a variable lets cancel y first -2(2x+3y)=-2(17) 3x+6y=30 -4x-6y=-34 3x+6y=30 COMBINE THESE TWO ____________ -x=-4 x=4 plug the x into one equation 2(4)+3y=17 8+3y=17 3y=9 y=3